# Parabolic induction in the homotopy category of pro-$p$ Iwahori-Hecke modules

### v2

@@ 4.2. Bousfield localisation
We note that (co)local objects and equivalences are usually defined in terms of homotopy function complexes in the literature (c.f. [13, Definition 3.1.4]). The above definition is equivalent to that because for pointed model categories the homotopy groups of these function complexes are precisely the ${\operatorname{Hom}}$ sets we considered above, see [14, Lemma 6.1.2]. With this notion at hand, we have the following definition of Bousfield localisation.
Definition 4.14 ([13, Definition 3.3.1]).
Let $\mathscr{M}$ be a pointed model category and $\mathrm{S}$ be a class of morphisms in $\mathscr{M}$. The right Bousfield localisation of $\mathscr{M}$ with respect to $\mathrm{S}$ is, if it exists, the model structure $\mathrm{R}_{\mathrm{S}}\mathscr{M}$ on the same underlying category as $\mathscr{M}$, with fibrations the same as in $\mathscr{M}$, weak equivalences the $\mathrm{S}$-colocal equivalences, and cofibrations the maps satisfying the left lifting property with respect to the trivial fibrations. The notion of left Bousfield localisation of $\mathscr{M}$ with respect to $\mathrm{S}$ is defined dually.
@@
    If $f$ is a fibration between fibrant objects, then $X$ is $f$-colocal if and only if
@@
These statements are proved in steps (1)-(3) of the proof of [7, Proposition 1.5.3] for injective model structure and the proof carries through mutatis mutandis in our more general setting. ∎
@@
    Suppose that the inclusion ${\mathcal{S}}\subset{\operatorname{Ho}}({\mathcal{A}})$ has a left adjoint $\mathbf{L}{\mathsf{L}}$. Let ${\mathcal{C}}_{\mathcal{S}}$ denote the class of objects of ${\mathcal{A}}$ whose image in ${\operatorname{Ho}}({\mathcal{A}})$ lies in ${\operatorname{Ker}}(\mathbf{L}{\mathsf{L}})$, and assume that $({\mathcal{C}}_{\mathcal{S}}\cap{\mathcal{C}},{\mathcal{A}},{\mathcal{W}}_{\mathcal{S}})$ defines an abelian model structure on ${\mathcal{A}}$. Then this model structure is the right Bousfield localisation of ${\mathcal{A}}$ with respect to $\mathrm{S}:=\{0\to X\mid X\in{\mathcal{W}}_{\mathcal{S}}\}$; and
@@
    Suppose that the inclusion ${\mathcal{S}}\subset{\operatorname{Ho}}({\mathcal{A}})$ has a right adjoint $\mathbf{R}{\mathsf{R}}$. Let ${\mathcal{F}}_{\mathcal{S}}$ denote the class of objects of ${\mathcal{A}}$ whose image in ${\operatorname{Ho}}({\mathcal{A}})$ lies in ${\operatorname{Ker}}(\mathbf{R}{\mathsf{R}})$, and assume that $({\mathcal{C}}\cap{\mathcal{W}}_{\mathcal{S}},{\mathcal{A}},{\mathcal{F}}_{\mathcal{S}})$ defines an abelian model structure on ${\mathcal{A}}$. Then this model structure is the right Bousfield localisation of ${\mathcal{A}}$ with respect to $\mathrm{S}^{\prime}:=\{0\to X\mid X\in{\mathcal{F}}_{\mathcal{S}}\}$.
@@
We only prove (i), the proof of (ii) being entirely analogous. By the dual of [20, Proposition 4.9.1], we have ${}^{\ddagger}{\mathcal{S}}={\operatorname{Ker}}(\mathbf{L}{\mathsf{L}})$ and ${\mathcal{S}}={\operatorname{Ker}}(\mathbf{L}{\mathsf{L}})^{\ddagger}$. By definition of $\mathrm{S}$, the class of $\mathrm{S}$-colocal objects is that of those cofibrant objects $X$ whose image in ${\operatorname{Ho}}({\mathcal{A}})$ lies in ${}^{\ddagger}{\mathcal{S}}={\operatorname{Ker}}(\mathbf{L}{\mathsf{L}})$, i.e. the class of $\mathrm{S}$-colocal objects is ${\mathcal{C}}_{\mathcal{S}}\cap{\mathcal{C}}$. Then 4.15 implies that an $\mathrm{S}$-colocal equivalence is a morphism which, in the arrow category of ${\operatorname{Ho}}({\mathcal{A}})$, is isomorphic to a epimorphism with kernel in ${\mathcal{W}}_{\mathcal{S}}$.
Now assume that $({\mathcal{C}}_{\mathcal{S}}\cap{\mathcal{C}},{\mathcal{A}},{\mathcal{W}}_{\mathcal{S}})$ defines an abelian model structure, which we denote by ${\mathcal{A}}_{\mathcal{S}}$. Then we must show that the weak equivalences in ${\mathcal{A}}_{\mathcal{S}}$ are the $\mathrm{S}$-colocal equivalences. For this, note first that ${\operatorname{id}}:{\mathcal{A}}\to{\mathcal{A}}_{\mathcal{S}}$ is right Quillen since ${\mathcal{W}}\subseteq{\mathcal{W}}_{\mathcal{S}}$, and we get a functor ${\operatorname{Ho}}({\operatorname{id}}):{\operatorname{Ho}}({\mathcal{A}})\to{\operatorname{Ho}}({\mathcal{A}}_{\mathcal{S}})$ by 2.8. Hence one gets that an $\mathrm{S}$-colocal equivalence is also isomorphic to a epimorphism with kernel in ${\mathcal{W}}_{\mathcal{S}}$ in the arrow category of ${\operatorname{Ho}}({\mathcal{A}}_{\mathcal{S}})$, and hence is a weak equivalence in ${\mathcal{A}}_{\mathcal{S}}$. For the converse, suppose that $f$ is a weak equivalence in ${\mathcal{A}}_{\mathcal{S}}$. Note first that ${\mathcal{C}}_{\mathcal{S}}\cap{\mathcal{C}}\cap{\mathcal{W}}_{\mathcal{S}}=\mathrm{Proj}({\mathcal{A}})={\mathcal{C}}\cap{\mathcal{W}}$, so that the trivial cofibrations in ${\mathcal{A}}$ and in ${\mathcal{A}}_{\mathcal{S}}$ coincide. By factorising $f$ in ${\mathcal{A}}_{\mathcal{S}}$ as a composite of a trivial cofibration and a trivial fibration, we therefore see that $f$ is isomorphic in the arrow category of ${\operatorname{Ho}}({\mathcal{A}})$ to a trivial fibration in ${\mathcal{A}}_{\mathcal{S}}$, i.e. an epimorphism with kernel in ${\mathcal{W}}_{\mathcal{S}}$, as required. ∎
@@ 5. Simple supersingular Hecke modules in the homotopy category
For the remainder of this paper, we assume that $k$ is algebraically closed. Our main goal is to study the isomorphism classes of simple ${\mathcal{H}}$-modules in ${\operatorname{Ho}}({\mathcal{H}})$. In light of 3.28(ii), we focus our attention on the case of simple supersingular modules.
@@ 5.1. The affine Hecke algebra and simple supersingular ${\mathcal{H}}$-modules
Fix a set $\{\hat{s}\mid s\in S\}\subseteq\widetilde{W}_{\operatorname{aff}}$ of lifts of the simple affine reflections defined as in [25, §4.8]. Then ${\mathcal{H}_{\operatorname{aff}}}$ is generated as a $k$-algebra by $\{T_{t}\mid t\in T(\mathbb{F}_{q})\}$ and $\{T_{\hat{s}}\mid s\in S\}$. In addition, each irreducible component $\Phi^{\prime}$ of $\Phi$ gives rise to a subset $S^{\prime}\subseteq S$ and a subalgebra of ${\mathcal{H}_{\operatorname{aff}}}$ generated by $k[T(\mathbb{F}_{q})]$ and $\{T_{\hat{s}}\mid s\in S^{\prime}\}$. We will refer to these subalgebras as the irreducible components of ${\mathcal{H}_{\operatorname{aff}}}$. We will also need a notion of irreducible components of the finite Hecke algebra ${\mathcal{H}}_{F}$ associated to a face $F\subseteq\overline{C}$. First, since $S$ is in bijection with the nodes of the affine Dynkin diagram of $G$, we may associate to the set $S_{F}$ the subgraph of the affine Dynkin diagram spanned by the corresponding nodes. Each connected component of that subgraph corresponds under the aforementioned bijection to a subset $S_{F^{\prime}}$ of $S_{F}$, itself corresponding to a face $F^{\prime}$ such that $F\subseteq\overline{F^{\prime}}$ by (2.3). We will call the associated subalgebras ${\mathcal{H}}_{F^{\prime}}$ of ${\mathcal{H}}_{F}$ arising in this way the irreducible components of ${\mathcal{H}}_{F}$.
Next, we recall the Ollivier-Schneider resolutions for ${\mathcal{H}}$-modules. For each $0\leq i\leq d=r_{ss}$, we write $\mathscr{F}_{i}$ for the set of $i$-dimensional faces $F$ of $\mathscr{X}$ which are contained in $\overline{C}$. Note that by [18, Lemma 1.2] this is a set of representatives for the $G_{\operatorname{aff}}$-orbits of the $i$-dimensional faces in $\mathscr{X}$. We also fix for each $i$ a set $\mathscr{F}_{(i)}\subseteq\mathscr{F}_{i}$ of representatives for the $G$-orbits of the $i$-dimensional faces in $\mathscr{X}$. Then [25, Theorem 3.14] implies that for every ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}})$ there is an exact sequence of ${\mathcal{H}}$-modules of the form
@@
where we note that our Hecke algebras are opposite to the ones in loc. cit. and hence our tensors products are reversed. Here, for a face $F\subseteq\overline{C}$, the twist of an ${\mathcal{H}}^{\dagger}_{F}$-module by the orientation character $\epsilon_{F}$ is the one defined in loc. cit., §3.3.1. We now see that there is a similar resolution for ${\mathcal{H}_{\operatorname{aff}}}$-modules, c.f. [19, Remark 4.1].
@@
The proof is completely analogous to the proof of [25, Theorem 3.14], working with $G_{\operatorname{aff}}$ instead of $G$, and is omitted. The methods of loc. cit., §1, then show that ${\mathcal{H}_{\operatorname{aff}}}$ is $n$-Gorenstein for some $n\leq d$. Thus the category ${\operatorname{Mod}}({\mathcal{H}_{\operatorname{aff}}})$ has a Gorenstein projective model structure and the homotopy category ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$ is well-defined. Furthermore, we obtain by tensoring with (5.2) an exact resolution
@@
Recall from [32, Proposition 2.2] that the characters of ${\mathcal{H}_{\operatorname{aff}}}$ are parametrized by pairs $(J,\xi)$ where $\xi$ is a character of $k[T(\mathbb{F}_{q})]$ and $J\subseteq S_{\xi}$ where $S_{\xi}$ is as in (2.2). The character $\chi$ corresponding to such a pair is defined by
@@
We will sometimes write $\chi=(J,\xi)$ to mean that $\chi$ is defined as above. When $S_{\xi}=S$, the character $\chi$ will be called a twisted trivial character, resp. twisted sign character, if it additionally satisfies $J=\emptyset$, resp. $J=S$. There is a notion of supersingularity for characters of ${\mathcal{H}_{\operatorname{aff}}}$, and we have that a character $\chi$ is supersingular if and only if its restriction to each irreducible component of ${\mathcal{H}_{\operatorname{aff}}}$ does not equal the restriction of a twisted sign or twisted trivial character (c.f. Theorem 6.15 in loc. cit). We will use this result as a definition of supersingularity for characters of ${\mathcal{H}_{\operatorname{aff}}}$. By convention, if $G=T$ then every character as above is supersingular.
Next, the group $\widetilde{\Omega}$ acts on the characters of ${\mathcal{H}_{\operatorname{aff}}}$ as follows. Given a character $\chi$ of ${\mathcal{H}_{\operatorname{aff}}}$ and $\tilde{\omega}\in\widetilde{\Omega}$, we let $\chi^{\tilde{\omega}}(h):=\chi(T_{\tilde{\omega}}hT_{\tilde{\omega}^{-1}})$ for $h\in{\mathcal{H}_{\operatorname{aff}}}$. We denote by $\widetilde{\Omega}_{\chi}$ the stabiliser of $\chi$ under this action. This is a subgroup of finite index in $\widetilde{\Omega}$, containing $T(\mathbb{F}_{q})$. We also let ${\mathcal{H}}_{\chi}$ denote the subalgebra of ${\mathcal{H}}$ generated by ${\mathcal{H}_{\operatorname{aff}}}$ and $\{T_{\tilde{\omega}}\mid\tilde{\omega}\in\widetilde{\Omega}_{\chi}\}$. It follows from the braid relations ${\mathcal{H}}_{\chi}$ is free over ${\mathcal{H}_{\operatorname{aff}}}$ and that ${\mathcal{H}}$ is free over ${\mathcal{H}}_{\chi}$, both as left and right modules. For future use, we record here the following:

### v4

@@ 4.2. Bousfield localisation
We note that (co)local objects and equivalences are usually defined in terms of homotopy function complexes in the literature (cf. [15, Definition 3.1.4]). The above definition is equivalent to that because for pointed model categories the homotopy groups of these function complexes are precisely the ${\operatorname{Hom}}$ sets we considered above, see [16, Lemma 6.1.2]. With this notion at hand, we have the following definition of Bousfield localisation.
Definition 4.14 ([15, Definition 3.3.1]).
Let $\mathscr{M}$ be a pointed model category and $\mathrm{S}$ be a class of morphisms in $\mathscr{M}$. The right Bousfield localisation of $\mathscr{M}$ with respect to $\mathrm{S}$ is, if it exists, the model structure $\mathscr{R}_{\mathrm{S}}\mathscr{M}$ on the same underlying category as $\mathscr{M}$, with fibrations the same as in $\mathscr{M}$, weak equivalences the $\mathrm{S}$-colocal equivalences, and cofibrations the maps satisfying the left lifting property with respect to the trivial fibrations. The notion of left Bousfield localisation of $\mathscr{M}$ with respect to $\mathrm{S}$ is defined dually.
@@
    If $f$ is a fibration between fibrant objects, then $X$ is $\{f\}$-colocal if and only if
@@
These statements are proved in steps (1)-(3) of the proof of [7, Proposition 1.5.3] for injective model structures and the proof carries through mutatis mutandis in our more general setting. ∎
@@
    Suppose that the inclusion ${\mathcal{S}}\subset{\operatorname{Ho}}({\mathcal{A}})$ has a left adjoint ${\mathsf{L}}_{{\mathcal{S}}}$. Let ${\mathcal{C}}_{\mathcal{S}}$ denote the class of objects of ${\mathcal{A}}$ whose image in ${\operatorname{Ho}}({\mathcal{A}})$ lies in ${\operatorname{Ker}}({\mathsf{L}}_{{\mathcal{S}}})$, and assume that $({\mathcal{C}}_{\mathcal{S}}\cap{\mathcal{C}},{\mathcal{A}},{\mathcal{W}}_{\mathcal{S}})$ defines an abelian model structure on ${\mathcal{A}}$. Then this model structure is the right Bousfield localisation of ${\mathcal{A}}$ with respect to $\mathrm{S}:=\{0\to X\mid X\in{\mathcal{W}}_{\mathcal{S}}\}$; and
@@
    Suppose that the inclusion ${\mathcal{S}}\subset{\operatorname{Ho}}({\mathcal{A}})$ has a right adjoint ${\mathsf{R}}_{{\mathcal{S}}}$. Let ${\mathcal{F}}_{\mathcal{S}}$ denote the class of objects of ${\mathcal{A}}$ whose image in ${\operatorname{Ho}}({\mathcal{A}})$ lies in ${\operatorname{Ker}}({\mathsf{R}}_{{\mathcal{S}}})$, and assume that $({\mathcal{C}}\cap{\mathcal{W}}_{\mathcal{S}},{\mathcal{A}},{\mathcal{F}}_{\mathcal{S}})$ defines an abelian model structure on ${\mathcal{A}}$. Then this model structure is the right Bousfield localisation of ${\mathcal{A}}$ with respect to $\mathrm{S}^{\prime}:=\{0\to X\mid X\in{\mathcal{F}}_{\mathcal{S}}\}$.
@@
We only prove (i), the proof of (ii) being entirely analogous. By the dual of 3.10(i), we have ${}^{\ddagger}{\mathcal{S}}={\operatorname{Ker}}({\mathsf{L}}_{{\mathcal{S}}})$ and ${\mathcal{S}}={\operatorname{Ker}}({\mathsf{L}}_{{\mathcal{S}}})^{\ddagger}$. By definition of $\mathrm{S}$, the class of $\mathrm{S}$-colocal objects is that of those cofibrant objects $X$ whose image in ${\operatorname{Ho}}({\mathcal{A}})$ lies in ${}^{\ddagger}{\mathcal{S}}={\operatorname{Ker}}({\mathsf{L}}_{{\mathcal{S}}})$, i.e. the class of $\mathrm{S}$-colocal objects is ${\mathcal{C}}_{\mathcal{S}}\cap{\mathcal{C}}$. Then 4.15 implies that an $\mathrm{S}$-colocal equivalence is a morphism which, in the arrow category of ${\operatorname{Ho}}({\mathcal{A}})$, is isomorphic to an epimorphism with kernel in ${\mathcal{W}}_{\mathcal{S}}$.
Now assume that $({\mathcal{C}}_{\mathcal{S}}\cap{\mathcal{C}},{\mathcal{A}},{\mathcal{W}}_{\mathcal{S}})$ defines an abelian model structure, which we denote by ${\mathcal{A}}_{\mathcal{S}}$. Then we must show that the weak equivalences in ${\mathcal{A}}_{\mathcal{S}}$ are the $\mathrm{S}$-colocal equivalences. For this, note first that ${\operatorname{id}}:{\mathcal{A}}\to{\mathcal{A}}_{\mathcal{S}}$ is right Quillen since ${\mathcal{W}}\subseteq{\mathcal{W}}_{\mathcal{S}}$, and we get a functor ${\operatorname{Ho}}({\operatorname{id}}):{\operatorname{Ho}}({\mathcal{A}})\to{\operatorname{Ho}}({\mathcal{A}}_{\mathcal{S}})$ by 2.8. Hence one gets that an $\mathrm{S}$-colocal equivalence is also isomorphic to an epimorphism with kernel in ${\mathcal{W}}_{\mathcal{S}}$ in the arrow category of ${\operatorname{Ho}}({\mathcal{A}}_{\mathcal{S}})$, and hence is a weak equivalence in ${\mathcal{A}}_{\mathcal{S}}$. For the converse, suppose that $f$ is a weak equivalence in ${\mathcal{A}}_{\mathcal{S}}$. Note first that ${\mathcal{C}}_{\mathcal{S}}\cap{\mathcal{C}}\cap{\mathcal{W}}_{\mathcal{S}}=\mathrm{Proj}({\mathcal{A}})={\mathcal{C}}\cap{\mathcal{W}}$, so that the trivial cofibrations in ${\mathcal{A}}$ and in ${\mathcal{A}}_{\mathcal{S}}$ coincide. By factorising $f$ in ${\mathcal{A}}_{\mathcal{S}}$ as a composite of a trivial cofibration and a trivial fibration, we therefore see that $f$ is isomorphic in the arrow category of ${\operatorname{Ho}}({\mathcal{A}})$ to a trivial fibration in ${\mathcal{A}}_{\mathcal{S}}$, i.e. an epimorphism with kernel in ${\mathcal{W}}_{\mathcal{S}}$, as required. ∎
@@ 5. Simple supersingular Hecke modules in the homotopy category
For the remainder of this paper, we assume that $k$ is algebraically closed. Our main goal is to study the isomorphism classes of simple ${\mathcal{H}}$-modules in ${\operatorname{Ho}}({\mathcal{H}})$. In light of 3.29(ii), we focus our attention on the case of simple supersingular modules.
@@ 5.1. The affine Hecke algebra and simple supersingular ${\mathcal{H}}$-modules
Recall that we fixed a set $\{\hat{s}\mid s\in S\}\subseteq\widetilde{W}_{\operatorname{aff}}$ of lifts of the simple affine reflections defined as in [27, §4.8]. Then ${\mathcal{H}_{\operatorname{aff}}}$ is generated as a $k$-algebra by $\{T_{t}\mid t\in T(\mathbb{F}_{q})\}$ and $\{T_{\hat{s}}\mid s\in S\}$. The ring structure of ${\mathcal{H}_{\operatorname{aff}}}$ is determined by the braid relations and the quadratic relations in (2.4). In addition, each irreducible component $\Phi^{\prime}$ of $\Phi$ gives rise to a subset $S^{\prime}\subseteq S$ and a subalgebra of ${\mathcal{H}_{\operatorname{aff}}}$ generated by $k[T(\mathbb{F}_{q})]$ and $\{T_{\hat{s}}\mid s\in S^{\prime}\}$. We will refer to these subalgebras as the irreducible components of ${\mathcal{H}_{\operatorname{aff}}}$. We will also need a notion of irreducible components of the finite Hecke algebra ${\mathcal{H}}_{F}$ associated to a face $F\subseteq\overline{C}$. First, since $S$ is in bijection with the nodes of the affine Dynkin diagram of $G$, we may associate to the set $S_{F}$ the subgraph of the affine Dynkin diagram spanned by the corresponding nodes. Each connected component of that subgraph corresponds under the aforementioned bijection to a subset $S_{F^{\prime}}$ of $S_{F}$, itself corresponding to a face $F^{\prime}$ such that $F\subseteq\overline{F^{\prime}}$ by (2.3). We call the associated subalgebras ${\mathcal{H}}_{F^{\prime}}$ of ${\mathcal{H}}_{F}$ arising in this way the irreducible components of ${\mathcal{H}}_{F}$.
Next, we recall the Ollivier-Schneider resolutions for ${\mathcal{H}}$-modules. For each $0\leq i\leq d=r_{ss}$, we write $\mathscr{F}_{i}$ for the set of $i$-dimensional faces $F$ of $\mathscr{X}$ which are contained in $\overline{C}$. Note that by [20, Lemma 1.2] this is a set of representatives for the $G_{\operatorname{aff}}$-orbits of the $i$-dimensional faces in $\mathscr{X}$. We also fix for each $i$ a set $\mathscr{F}_{(i)}\subseteq\mathscr{F}_{i}$ of representatives for the $G$-orbits of the $i$-dimensional faces in $\mathscr{X}$. Then [27, Theorem 3.12] implies that for every ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}})$ there is an exact sequence of ${\mathcal{H}}$-modules of the form
@@
where we note that our Hecke algebras are opposite to the ones in loc. cit. and hence our tensors products are reversed. Here, for a face $F\subseteq\overline{C}$, the twist $(\epsilon_{F}){\mathcal{H}}$ by the orientation character $\epsilon_{F}$ is the one defined in loc. cit., §3.3.1. We now see that there is a similar resolution for ${\mathcal{H}_{\operatorname{aff}}}$-modules, cf. [21, Remark 4.1].
@@
Proof.
The proof is mostly analogous to that of [27, Theorem 3.12], we only sketch how it goes through for ${\mathcal{H}_{\operatorname{aff}}}$ instead of ${\mathcal{H}}$. First, recall that to any $V\in{\operatorname{Rep}}_{k}^{\infty}(G)$, we may functorially associate a certain $G$-equivariant coefficient system $\underline{\underline{V}}$ as in [20, Example 2.2]. We refer the reader to §2.1 of loc. cit. for the relevant definitions, we only point out that this construction also applies, with the same definitions, to smooth representations of $G_{\operatorname{aff}}$ and that there is a functorial oriented chain complex ${\mathcal{C}}^{or}_{c}(\mathscr{X}_{(\bullet)},\underline{\underline{V}})\to V$ for $V$ a smooth representation of one of $G$, $G_{\operatorname{aff}}$ (cf. equations (2.1)-(2.2) of loc. cit.). If ${\mathbb{X}}$ and ${\mathbb{X}}_{\operatorname{aff}}$ are the representations from §2.2, then note that extending functions by zero gives a $G_{\operatorname{aff}}$-equivariant split embedding ${\mathbb{X}}_{\operatorname{aff}}\hookrightarrow{\mathbb{X}}$ (with splitting given by restricting functions to $G_{\operatorname{aff}}$). Thus we see that the complex $0\to{\mathcal{C}}^{or}_{c}(\mathscr{X}_{(\bullet)},\underline{\underline{{\mathbb{X}}_{\operatorname{aff}}}})\to{\mathbb{X}}_{\operatorname{aff}}\to 0$ is a $G_{\operatorname{aff}}$-equivariant retract of $0\to{\mathcal{C}}^{or}_{c}(\mathscr{X}_{(\bullet)},\underline{\underline{{\mathbb{X}}}})\to{\mathbb{X}}\to 0$. Since the latter is exact on $I$-invariants by [27, Theorem 3.4] it follows that the same holds for the former. We are therefore left to show that the ${\mathcal{H}_{\operatorname{aff}}}$-bimodule ${\mathcal{C}}^{or}_{c}(\mathscr{X}_{(i)},\underline{\underline{{\mathbb{X}}_{\operatorname{aff}}}})^{I}$ is as given in the statement for each $0\leq i\leq d$.
This is done by arguing exactly as in [27]. The proof that
${\mathcal{C}}^{or}_{c}(\mathscr{X}_{(i)},\underline{\underline{{\mathbb{X}}_{\operatorname{aff}}}})\cong\bigoplus_{F\in\mathscr{F}_{i}}{\operatorname{ind}}_{P_{F}}^{G_{\operatorname{aff}}}({\mathbb{X}}_{\operatorname{aff}}^{I_{F}})$
is done exactly as in §3.3.2 of loc. cit. We only point out that the orientation character $\epsilon_{F}$ doesn’t appear here because each parahoric $P_{F}$ fixes $F$ pointwise and thus also fixes any chosen orientation on $F$, and that $\mathscr{F}_{i}$ is a set of representatives for the $G_{\operatorname{aff}}$-orbits of the $i$-dimensional faces in $\mathscr{X}$. The proof of Proposition 4.25 in loc. cit. then adapts to give an isomorphism ${\mathcal{H}_{\operatorname{aff}}}\otimes_{{\mathcal{H}}_{F}}{\mathbb{X}}_{F}\cong{\mathbb{X}}_{\operatorname{aff}}^{I_{F}}$ of left $k[P_{F}]\otimes_{k}{\mathcal{H}_{\operatorname{aff}}}$-modules. Therefore, one obtains
${\operatorname{ind}}_{P_{F}}^{G_{\operatorname{aff}}}({\mathbb{X}}_{\operatorname{aff}}^{I_{F}})^{I}\cong{\operatorname{ind}}_{P_{F}}^{G_{\operatorname{aff}}}({\mathcal{H}_{\operatorname{aff}}}\otimes_{{\mathcal{H}}_{F}}{\mathbb{X}}_{F})^{I}\cong{\mathcal{H}_{\operatorname{aff}}}\otimes_{{\mathcal{H}}_{F}}{\mathcal{H}_{\operatorname{aff}}}$
exactly as in Proposition 3.10 of loc. cit. ∎
The methods of [27, §1] then show that ${\mathcal{H}_{\operatorname{aff}}}$ is $n$-Gorenstein for some $n\leq d$. Thus the category ${\operatorname{Mod}}({\mathcal{H}_{\operatorname{aff}}})$ has a Gorenstein projective model structure and the homotopy category ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$ is well-defined. Furthermore, we obtain by tensoring with (5.2) an exact resolution
@@
Recall from [35, Proposition 2.2] that the characters of ${\mathcal{H}_{\operatorname{aff}}}$ are parametrized by pairs $(J,\xi)$ where $\xi$ is a character of $k[T(\mathbb{F}_{q})]$ and $J\subseteq S_{\xi}$ where $S_{\xi}$ is as in (2.2). The character $\chi$ corresponding to such a pair is defined by
@@
We will sometimes write $\chi=(J,\xi)$ to mean that $\chi$ is defined as above. When $S_{\xi}=S$, the character $\chi$ will be called a twisted trivial character, resp. twisted sign character, if it additionally satisfies $J=\emptyset$, resp. $J=S$. There is a notion of supersingularity for characters of ${\mathcal{H}_{\operatorname{aff}}}$, and we have that a character $\chi$ is supersingular if and only if its restriction to each irreducible component of ${\mathcal{H}_{\operatorname{aff}}}$ does not equal the restriction of a twisted sign or twisted trivial character (cf. Theorem 6.15 in loc. cit). We will use this result as a definition of supersingularity for characters of ${\mathcal{H}_{\operatorname{aff}}}$. By convention, if $G=T$ then every character as above is supersingular.
Next, the group $\widetilde{\Omega}$ acts on the characters of ${\mathcal{H}_{\operatorname{aff}}}$ as follows. Given a character $\chi$ of ${\mathcal{H}_{\operatorname{aff}}}$ and $\tilde{\omega}\in\widetilde{\Omega}$, we let $\chi^{\tilde{\omega}}(h):=\chi(T_{\tilde{\omega}}hT_{\tilde{\omega}^{-1}})$ for $h\in{\mathcal{H}_{\operatorname{aff}}}$. We denote by $\widetilde{\Omega}_{\chi}$ the stabiliser of $\chi$ under this action. This is a subgroup of finite index in $\widetilde{\Omega}$, containing $T(\mathbb{F}_{q})$. We also let ${\mathcal{H}}_{\chi}$ denote the subalgebra of ${\mathcal{H}}$ generated by ${\mathcal{H}_{\operatorname{aff}}}$ and $\{T_{\tilde{\omega}}\mid\tilde{\omega}\in\widetilde{\Omega}_{\chi}\}$. It follows from the braid relations that ${\mathcal{H}}_{\chi}$ is free over ${\mathcal{H}_{\operatorname{aff}}}$ and that ${\mathcal{H}}$ is free over ${\mathcal{H}}_{\chi}$, both as left and right modules. For future use, we record here the following: